I'm not sure what category this problem falls under. In the problem 2x + 3y = 7 what is the expression for y in terms of x?
Answer by mananth(16946)

2x+3y=7
add -2x
2x-2x+3y=7-2x
3y=7-2x
divide by 3
3y/3 = (7-2x)/3
y= (7-2x)/3